OpenStudy (anonymous):

One number is three less than four times another. If their difference is 48, what is the larger number? a.17 b.51 c.65

OpenStudy (anonymous):

Let x = first number y = second number x = 4y-3 so for the equation we would be using y and (4y-3) we assume that 4y-3 is the larger number. their difference is (larger) - (smaller) = 48 (4y-3) - (y) = 48 4y - 3 - y = 48 3y = 51 y = 17 4y - 3 = 4(17) - 3 = 68 -3 = 65